Ticket Number: HEA-NOC/20051112-3 Ticket Status: OPEN
Ticket Type: unscheduled Resolver: HEA
Ticket Opened: 20051112 12:57 UTC Problem Start: 20051112 11:29 UTC
Ticket Update: 20051112 12:57 UTC Problem End:
Our monitoring tools are reporting alarms for the HEA since 11:29. Link is
local copper and our router interface reports up/up, so a problem with the
remote equipment is suspected.
All traffic to and from the HEA is affected.
20051112 12:55 DW Left message for HEA technical contact with out of hours
Time to Fix:
Not yet known.
This ticket can be monitored at http://www.hea.net/tickets/20051112-3